A kind teacher gives you 20 cents for every question you get right, but you have to pay the teacher 10 cents for every question you get wrong. After 30 questions you have made a profit of $1.80.
a) Form an equation with X representing the number of questions you got right
B) Solve your equation to find how many questions you got right.

Found 2 solutions by ikleyn, solver91311:
Answer by ikleyn(52908)

.

(a)  This equation is

         20*X - 10*(30-X) = 180  cents.


(b)  From the equation

         20X - 300 + 10X = 180

         30X             = 180 + 300

         30X             = 480

           X             = 480/30 = 16.


ANSWER.  You got right 16 questions.


CHECK.   20*16 - 10*(30-16) = 180 cents = $1.80, your profit (!).   Precisely correct (!)

Solved.



Answer by solver91311(24713)



Let represent the number of questions answered correctly, and let represent the number of questions answered incorrectly. Then:





Note $1.80 is 180 cents

Solve the 2X2 system for and


John

My calculator said it, I believe it, that settles it